You have three electrical appliances (1) 1KW-250 V electric heater , (2) 1KW-250 V electric kettle and (3)1KW-259 V electric bulb.Which of these have highest resistance?
1
Expert's answer
2012-07-05T09:44:24-0400
P = UI = U^2 / R

As P1 = P2 = P3 = 1KW, the electric bulb has highest resistance (U3 > U2 = U1).

R3 = U^2 / P = 259^2 / 1000 = 67.081 Ohm.